Calculate the Qsp for the mixing of 30.0mL of 3.7 × 10^-3 M Sr(NO3)2 and 25.0mL of 0.02 M Na2SO4


1
Expert's answer
2021-01-14T04:25:29-0500

Qsp of SrSO4 =

Concentration of Sr2+ × Concentration of SO42- .


= ( 30 × 3.7 × 10-6 ) / 55 × 10-3

×

( 25 × 0.02 × 10-6 ) / 55 × 10-3 ).


= 30 × 3.7 × 25 × 0.02 × 10-6/ 55 × 55


= 1.8 × 10-8 ..